# Harmony: A Joint Self-Supervised and Weakly-Supervised Framework for Learning General Purpose Visual Representations

### v1

@@ 4.2 Results
Size R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
CLIP [40] S 3.9 11.7 17.5 3.9 10.9 15.8 13.1 32.8 44.4 11.7 29.6 40.4
SLIP [37] S 5.4 14.9 21.9 4.5 12.3 17.5 16.3 36.6 49.2 12.6 30.3 41.2
MaskCLIP [17] S 4.5 12.6 18.4 4.3 11.7 16.9 13.4 33.5 45.7 12.2 30.5 41.0
Harmony S 8.1 20.5 28.7 7.4 18.2 25.0 19.8 42.8 55.8 17.0 38.8 50.3
Harmony B 9.7 24.0 32.9 8.7 20.9 28.4 23.0 46.6 58.6 19.1 42.2 54.0
@@ 4.3 Ablations
Building Harmony. To quantify the improvements from each objective in Harmony, we rebuild Harmony’s architecture step-by-step, starting at CLIP [40]. The results for pre-training a CLIP ViT-S/16 on CC3M [43] for 25 epochs are shown in Table 4. We continue by adding a visual feature prediction task in the form of self-distillation. This substantially boosts the accuracies across all tasks. Subsequently, we add soft targets by utilizing the same teacher network $\bar{E}_{V}$ used in the feature prediction, further increasing performance. However, we notice a degradation in the SSEG performance. To remedy this, we add the pixel reconstruction loss from MAE [24], which operates in the pixel space, leading to more granular pixel-level features being learned. This increases SSEG closer its value after feature prediction. The 0-shot and LIN accuracies also increase slightly.
@@
Moreover, we investigate the effect of adding our two text losses to the original CLIP. Starting from the same baseline show in Table 4, adding both MLM and text self-distillation objectives increases the 0-shot classification by +1.3% compared to CLIP. This is likely due to a boost in representation ability of the text encoder ${E}_{T}$, and the introduction of an EMA teacher text encoder $\bar{E}_{T}$. However, even in downstream tasks that do not utilize text encoders like FT and SSEG, there is still a slight increase in performance, likely due to the making the contrastive learning task more meaningful by enhancing textual representations, leading to more reflective similarity matrices in Equation 3.
Table 4: Ablation study for Harmony. The upper section represents the addition of the main three objectives of our framework. The feature prediction task is described in Section 3.2, soft targets in Section 3.1 and pixel reconstruction in Section 3.3. The lower part shows the addition of the two text losses. See Appendix section B for more information about compute. We use a batch size of 512 and ViT-S/16 for all our ablations. Text losses and Harmony are compared to the original CLIP baseline.
Method Compute INET-1K (ACC) ADE20K (mIoU)
Mem Time 0-Shot FT LIN SSEG
CLIP 1.0x 1.0x 16.7 76.8 56.4 35.3
+ Feature Prediction 2.4x 3.0x 19.2 ↑2.5 78.3 ↑1.5 61.5 ↑5.1 36.5 ↑1.2
+ Soft Targets 2.7x 3.1x 20.5 ↑1.3 79.5 ↑1.2 64.2 ↑2.7 35.3 ↓1.2
+ Pixel Reconstruction 3.5x 3.7x 20.8 ↑0.3 79.5 ↑0.0 64.4 ↑0.2 36.2 ↑0.9
Text losses
+ MLM 1.6x 1.6x 17.3 ↑0.6 77.1 ↑0.3 35.4 ↑0.1
+ Text Self-distillation 1.6x 1.7x 18.0 ↑0.7 77.1 ↑0.0 35.6 ↑0.2
= Harmony 4.1x 4.1x 21.0 ↑4.3 79.8 ↑3.0 64.6 ↑8.2 37.9 ↑2.6
@@
Hyper-parameter tuning. We focus mainly on the introduced $\alpha_{c}$ parameter of Equation 3.1 that controls the weight of soft and hard CLIP targets, and the parameters in Equation 13 that control the influence of each objective in Harmony. In Appendix D, we provide additional experiments such as the effect of passing the two global crops compared to a minimally augmented image to our pixel reconstruction objective, using iBOT’s data augmentation with CLIP, using an mask scheduler to change the MAE mask ratio throughout training, and masking the images passed to CLIP [31, 51].
@@
Influence of $\alpha_{c}$ for soft targets. For the $\alpha_{c}$ parameter in the contrastive loss, we compare two situations: a higher alpha value ($\alpha_{c}$ = 0.2) reached early in the training (10 epochs), compared to a lower alpha ($\alpha_{c}=0.05$), reached later in the training (15 epochs). We show both results on the left part of Table 5. A higher value performs significantly better.
@@
Loss weight. In the right part of Table 5, we adjust the weight parameter for each of the four loss functions added on top of the contrastive loss. We observe slight performance changes, with setting the weight to one generally performing better.
@@
Table 5: ImageNet-1k 0-shot for different hyper-parameters. Left: Alpha value for soft labels. Higher values works better. Both start with same baseline of CLIP + iBOT. Right: Loss Weighting. We observe that setting the weight to one performs slightly better. We compare CLIP + iBOT for $\alpha$, CLIP + MAE for $\beta$, CLIP + MLM for $\gamma$, and CLIP + MLM + TextDist for $\delta$ with $\gamma$ set to 1.
Alpha $\alpha_{c}$ 1 $\rightarrow$ 0.2 in 10 epochs 1 $\rightarrow$ 0.05 in 15 epochs
0-shot 20.5 16.6
@@ 5 Conclusion
We introduce Harmony, a joint self-supervised and weakly-supervised method for learning generalized visual features from web-scraped data. Harmony outperforms previous methods and baselines across classification, segmentation, and detection tasks, highlighting how our multiple training objectives can complement each other to learn stronger visual representations.
Acknowledgements
@@ Appendix A Experimental Details
Pre-training. We use the AdamW optimizer [33], and mixed precision with FP16, in all our pre-training experiments. Following DINO [6] and iBOT [55] we scale the learning rate with the batch size through $lr=5e^{-4}\times batchsize/256$. The learning rate is ramped up linearly from 0 to the value in the formula in the first 3 epochs, then decreases to $1e^{-6}$ using a cosine scheduler, by the end of training. For iBOT and DINO heads, we set the output dimension to 8192, and we share weights in both the iBOT and DINO heads for both the student and teacher. We also normalize the last layer for stability, and set the momentum parameter for the teacher to 0.996. For the masked image modeling in iBOT [55], we follow the settings in their ViT-B/16 ImageNet-22k experiments. Specifically, the prediction ratio is selected randomly as either 0 and 0.3 plus an added variation of ±$\pm 0.2$. Masking is performed block-wise. For the CLS objective in Equation 6, we set the temperature parameter $\tau$ to 0.04. For the MLM objective shown in Equation 8, the temperature parameter is linearly scaled from 0.04 to 0.07 in 10 epochs. We use a weight decay that is scaled with a cosine scheduler from 0.04 to 0.4, and use gradient clipping with a max gradient norm value of 3.0.
@@
For data augmentation, we likewise follow the settings in iBOT [55]. For the global crops, we use a cropping with a scale uniformly sampled from $\left[0.32,1\right]$. In all our ViT-S/16 experiments, we use 8 local crops in the global objective, while we use 6 local crops for ViT-B/16 experiments, all with cropping scale sampled from $\left[0.05,0.32\right]$. For the first global crop, we always apply a Gaussian blur, while we randomly apply it to the second crop with a probability of 0.1. We also randomly solarize the second crop with a probability of 0.2 and threshold of 128. We also apply a horizontal flipping to both crops with a probability of 0.5, a color jitter with probability of 0.8, and grayscaling with probability 0.2. For the local crops, we take random crops of size 96 $\times$ 96, and apply the same horizontal flipping and color jitter. We also apply a Gaussian blur with probability 0.5.
@@
For the MAE pixel reconstruction objective, we randomly mask (remove) 75% of all patch tokens following [24]. We reconstruct both global crops from iBOT (see 10). The final pixel reconstruction loss is then calculated as the addition of the loss of both global crops. For the MLM, objective, each token is randomly masked with 20% probability.
For all ViT-S/16 experiments, we use between 8 and 16 V100 GPUS running for a day. For the ViT-B/16 run, we use 24 GPUs insted, running for a day.
@@
Fine-tuning on ImageNet-1k. For fine-tuning evaluations, we use the pipeline in iBOT [55], which follows the fine-tuning protocol from BEiT [3]. We fine-tune for 100 epochs in all experiments using the AdamW [33] optimizer, with a batch size of 1024. We linearly warmup the learning rate to $4e^{-3}$ in 20 epochs then lower it with a cosine scheulder to a final value of $1e^{-6}$. We use a layer-wise decay of 0.65, and a weight decay of 0.05. We do not use layer scale. We use 8 32GB V100 GPUs for our fine-tuning experiments.
@@
Linear-probing classification on ImageNet-1k. For our linear-probing experiments, we mostly follow DINO [6] and DINOv2 [38]. We train for 100 epochs and a batch size of 8192 in all our experiments. We use the CLS token only from the last layer of the ViT encoder. Following [38], we instantiate multiple linear layers that all take the same encoder output, but are trained with different learning rates for efficiency. This way, the larger backbone inference is done once per iteration, while multiple inferences are done on the much lighter linear layers. We specifically span learning rate values of $\{1e^{-5},2e^{-5},5e^{-5},1e^{-4},2e^{-4},5e^{-4},1e^{-3},2e^{-3},5e^{-3},1e^{-2},2e^{-2},5e^{-2},1e^{-1}\}$, and we linearly scale them using the formula $lr\times batchsize/256$. We report the best performing value. We use a stochastic gradient descent optimizer with the momentum set to 0.9 and no weight decay. We use 8 32GB V100 GPUs for our linear-probing experiments.
@@
ADE20K semantic segmentation. In our semantic segmentation evaluations, we also use the iBOT [55] pipeline, which follows BEiT [3]. Specifically, we use UperNet [49] from the implementation in mmsegmentaion [14]. We fine-tune for 160k iterations using a batch size of 16 and image size of 512 $\times$ 512. The the AdamW optimizer is used with an initial learning rate of $8e^{-4}$ that is linearly warmed in the first 1500 epochs, then decays to 0 throughout training. We use a layer decay rate of 0.65 and weight decay of 0.05. We use Feature Pyramid Networks (FPNs) of four different scales to modify the feature map sizes generated by the ViT. Specifically, we upsample the output feature of the 4th block and 6th block by 4x and 2×, respectively, keep the output from the 8th block the same, and downsample the output feature of the 12th block by 2×. For data augmentation, we adopt the default settings in mmsegmentation, which includes random horizontal flipping, random re-scaling with a ratio range of $\left[0.5,2.0\right]$, and random photometric distortion. We use 4 32GB V100 GPUs for our semantic segmentation experiments
@@
COCO object detection and instance segmentation. For object detection and instance segmentation, we use a Cascade Mask R-CNN [4, 22] implementation with mmdetection [8], which produces both bounding boxes and instance masks. We follow the settings in iBOT [55] in that we use multi-scale training with the shorter size between 480 and 800 while the longer size is not larger than 1333. We use a learning rate of $1e^{-4}$, a weight decay of 0.05, and fine-tune for 12 epochs with the learning rate decayed with a rate of 0.1 at epochs 9 and 11. We use a layer decay rate of 0.75 and a batch size of 16. We generate hierarchical feature maps by taking the outputs from layers 4, 6, 8, and 12, and passing them to two deconvolutions, one deconvolution, identity mapping, and max-pooling, respectively. We do not use multi-scale testing. We use 4 32GB V100 GPUs for our object detection experiments.
Appendix B Memory Usage and Training Efficiency
Since Harmony utilizes iBOT [55], which defines two global crops per image, and multiple local crops, the relative difference in compute between Harmony and iBOT keeps increasing as we increase the batch size. We show values for different batch sizes in Table 6. Here, memory is measured as the maximum GPU memory usage reached in an epoch, and time is calculated as the time required to complete one epoch, and we use 10k images per epoch. When the batch size is set to 1, most of the memory differences come from additional model parameters, as opposed to images in a batch. As we scale the batch size, compute increases at a slower rate, since the percentage of compute taken by the model parameters becomes smaller, and differences in compute become easier to model as differences in images per batch size.
Table 6: Compute efficiency of Harmony compared to the CLIP baseline at same batch sizes. For example, for a batch size of 8, Harmony requires 3 times the memory as CLIP.
Per GPU Batch Size Memory Time
1 2.0x 3.2x
8 3.0x 3.6x
16 3.5x 3.9x
32 4.1x 4.1x

### v3

@@ 4.2 Results
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
CLIP 4.9 13.8 20.2 4.6 12.6 18.1 15.4 36.2 48.0 13.8 32.9 43.9
MaskCLIP 4.8 14.0 20.5 4.7 12.5 17.9 16.1 36.3 48.1 13.4 32.0 42.8
DetailCLIP 6.4 16.8 24.2 5.8 14.8 20.7 17.9 39.4 51.3 14.8 33.7 44.8
SLIP 5.4 15.4 22.3 5.0 13.5 19.1 15.4 38.0 50.3 13.4 32.3 43.2
SILC 7.6 20.4 28.9 7.3 18.4 25.4 20.6 44.5 56.5 17.9 39.7 51.4
Harmony 9.8 23.7 32.6 8.6 20.7 28.1 22.1 46.9 58.7 19.1 42.1 53.8
Table 4: Ablation study for Harmony. The upper section represents the addition of the main three objectives of our framework. The feature prediction task ($\mathcal{L}_{D}$) is described in Section 3.2, soft targets ($\mathcal{L}_{\text{Soft}}$) in Section 3.1 and pixel reconstruction ($\mathcal{L}_{R}$) in Section 3.3. The lower part shows the addition of the two text losses: Masked language modeling ($\mathcal{L}_{M}$) and text self-distillation ($\mathcal{L}_{TD}$). We train a ViT-S for 25 epochs and a batch size 768 for all ablations. Text losses are compared to the original CLIP baseline.
Method Compute INet-1K ADE20K
Mem Time 0-Shot FT LIN SSEG
CLIP 1.0x 1.0x 17.2 76.7 57.0 36.5
+ $\mathcal{L}_{D}$ 2.3x 3.0x 19.8 ↑2.6 78.2 ↑1.5 61.7 ↑4.7 37.9 ↑1.4
+ $\mathcal{L}_{\text{Soft}}$ 2.3x 3.2x 20.4 ↑0.6 79.4 ↑1.2 64.3 ↑2.6 37.0 ↓0.9
+ $\mathcal{L}_{R}$ 2.4x 3.9x 20.4 ↑0.0 79.5 ↑0.1 64.4 ↑0.1 38.0 ↑1.0
Text
+ $\mathcal{L}_{M}$ 1.1x 1.3x 17.7 ↑0.5 76.8 ↑0.1 35.8 ↓0.7
+ $\mathcal{L}_{TD}$ 1.1x 1.4x 18.6 ↑0.9 77.1 ↑0.3 36.4 ↑0.6
= $\mathcal{L}_{H}$ 2.5x 4.2x 20.8 ↑3.6 79.5 ↑2.8 64.4 ↑7.4 37.5 ↑1.0
@@ 4.3 Ablations
Building Harmony. To quantify the improvements from each objective in Harmony, we rebuild Harmony’s architecture step-by-step, starting at CLIP (Radford et al., 2021). The results for pre-training a CLIP on CC3M (Sharma et al., 2018) are shown in Table 4. We continue by adding a visual feature prediction task in the form of self-distillation ($\mathcal{L}_{D}$). This substantially boosts the accuracies across all tasks. Subsequently, we add soft targets ($\mathcal{L}_{\text{Soft}}$) by utilizing the same teacher network $\bar{E}_{V}$ used in the feature prediction, further increasing performance. However, we notice a degradation in the SSEG performance. To remedy this, we add the pixel reconstruction loss ($\mathcal{L}_{R}$) from MAE (He et al., 2021), which operates in the pixel space, leading to more granular pixel-level features being learned, increasing SSEG closer its value after feature prediction.
@@
Moreover, we investigate the effect of adding our two text losses to the original CLIP. Starting from the same baseline show in Table 4, adding both MLM ($\mathcal{L}_{M}$) and text self-distillation ($\mathcal{L}_{TD}$) objectives increases the 0-shot classification by +1.4% compared to CLIP. This is likely due to a boost in representation ability of the text encoder ${E}_{T}$, and the introduction of an EMA teacher text encoder $\bar{E}_{T}$. However, even in downstream tasks that do not utilize text encoders like FT, there is still a slight increase in performance, likely due to the making the contrastive learning task more meaningful by enhancing textual representations, leading to more reflective similarity matrices in Equation 3.
@@
Hyper-parameter tuning. We tune the introduced $\alpha_{c}$ parameter of Equation 5 that controls the weight of soft and hard CLIP targets, and the parameters in Equation 13 that control the influence of each objective in Harmony. In Appendix D, we provide additional experiments such as the effect of (1) passing the two global crops compared to a single image to our pixel reconstruction objective, (2) using iBOT’s data augmentation with CLIP, (3) using a mask scheduler to change the MAE mask ratio throughout training, and (4) masking the images passed to CLIP (Li et al., 2023; Yang et al., 2023).
Table 5: Equal compute comparison. Performance of CLIP, SLIP, and MaskCLIP when trained with the same hyperparameters and the same computational resources (light gray) as Harmony. We saved a checkpoint every 10 epochs and reported results for the highest performance on zero-shot classification, because the models were overfitting due to the larger number of epochs. Total memory is calculated across all GPUs used, and GPU hour is calculated for a single GPU.
Method GPU Hour Total Memory (GB) Batch Size Epochs 0-Shot FT
CLIP 90 76 768 25 17.2 76.7
CLIP 413 209 2400 120 17.9 77.7
SLIP 167 184 768 25 18.1 76.5
SLIP 426 240 1024 65 18.4 77.7
MaskCLIP 105 197 768 25 17.0 77.0
MaskCLIP 417 202 800 100 19.6 78.4
Harmony 414 219 768 25 20.8 79.5
@@
Influence of $\alpha_{c}$ for soft targets. For the $\alpha_{c}$ parameter in the contrastive loss, we compare two situations: A higher alpha value ($\alpha_{c}$ = 0.2) reached early in the training (10 epochs) results in a 0-shot accuracy of 20.5, while a lower alpha ($\alpha_{c}=0.05$), reached later in the training (15 epochs) results in a 0-shot accuracy of 16.6.
@@
Loss weight. In Table 6, we adjust the weight parameter for each of the four loss functions added on top of the contrastive loss. We observe slight performance changes, with setting the weight to one generally performing better.
Batch size. We investigated the effect of changing the batch size for Harmony, using a ViT-S. Table 7 shows the results. The performance for 0-shot classification gradually increases as we go from a batch size of 512 to 1024, but then decreases when we go to 2048.
Equal compute. As shown in Table 4, Harmony uses 2.5 times more compute and takes 4.2 times more time to train for the same number of epochs, compared to CLIP, SLIP and MaskCLIP. This makes a direct comparison between the two methods, while controlling for the number of epochs is unintuitive. Instead, in Table 5, we control for memory and time rather than the number of epochs. We increase the batch size to use more memory for the other methods. Harmony still significantly outperforms the other methods, which seem to plateau after 25 epochs, increasing their accuracy only by 1% or less (except MaskCLIP). For each run, checkpoints were saved every 10 epochs, and only the best result is shown, because the models were overfitting due to the larger number of epochs.
@@
Table 6: ImageNet-1k 0-shot accuracy for different Loss Weighting. We observe that setting the weight to one performs slightly better. We compare CLIP + iBOT for $\alpha$, CLIP + MAE for $\beta$, CLIP + MLM for $\gamma$, and CLIP + MLM + TextDist for $\delta$ with $\gamma$ set to 1.
@@
Table 7: Influence of changing Harmony’s batch size.
Batch Size 512 768 1024 2048
0-Shot 20.3 20.8 21.1 20.6
@@ 5 Conclusion
We present Harmony, a joint self-supervised and weakly-supervised method for learning generalized visual features from web-scraped data, introducing a soft loss and a text self-distillation method. Harmony outperforms or is competitive with leading methods and baselines across classification, segmentation, and detection tasks, highlighting how our multiple training objectives can complement each other to learn stronger visual representations.
6 Acknowledgement
@@ Appendix A Experimental Details
We use the AdamW optimizer (Loshchilov & Hutter, 2019), and mixed precision with FP16, in all our pre-training experiments. Following DINO (Caron et al., 2021b) and iBOT (Zhou et al., 2022) we scale the learning rate with the batch size through $lr=5e^{-4}\times batchsize/256$. The learning rate is ramped up linearly from 0 to the value in the formula in the first 3 epochs, then decreases to $1e^{-6}$ using a cosine scheduler, by the end of training. For iBOT and DINO heads, we set the output dimension to 8192, and we share weights in both the iBOT and DINO heads for both the student and teacher. We also normalize the last layer for stability, and set the momentum parameter for the teacher to 0.996. For the masked image modeling in iBOT (Zhou et al., 2022), we follow the settings in their ViT-B ImageNet-22k experiments. Specifically, the prediction ratio is selected randomly as either 0 and 0.3 plus an added variation of ±$\pm 0.2$. Masking is performed block-wise. For the CLS objective in Equation 6, we set the temperature parameter $\tau$ to 0.04. For the MLM objective shown in Equation 8, the temperature parameter is linearly scaled from 0.04 to 0.07 in 10 epochs. We use a weight decay that is scaled with a cosine scheduler from 0.04 to 0.4, and use gradient clipping with a max gradient norm value of 3.0.
@@
For data augmentation, we likewise follow the settings in iBOT (Zhou et al., 2022). For the global crops, we use a cropping with a scale uniformly sampled from $\left[0.32,1\right]$. In all our experiments, we use 8 local crops in the global objective, with cropping scale sampled from $\left[0.05,0.32\right]$. For the first global crop, we always apply a Gaussian blur, while we randomly apply it to the second crop with a probability of 0.1. We also randomly solarize the second crop with a probability of 0.2 and threshold of 128. We also apply a horizontal flipping to both crops with a probability of 0.5, a color jitter with probability of 0.8, and grayscaling with probability 0.2. For the local crops, we take random crops of size 96 $\times$ 96, and apply the same horizontal flipping and color jitter. We also apply a Gaussian blur with probability 0.5.
@@
For the MAE pixel reconstruction objective, we randomly mask (remove) 75% of all patch tokens following He et al. (2021). We reconstruct both global crops from iBOT (see Table 12). The final pixel reconstruction loss is then calculated as the addition of the loss of both global crops. For the MLM, objective, each token is randomly masked with 20% probability.
For all ViT-S and ViT-B experiments, we use between 8 and 16 V100 GPUS running for a day. For the ViT-L run, we use 16 80GB A100 GPUs instead, running for a day.
@@
Fine-tuning on ImageNet-1k. For fine-tuning evaluations, we use the pipeline in iBOT (Zhou et al., 2022), which follows the fine-tuning protocol from BEiT (Bao et al., 2022). We fine-tune for 100 epochs in all experiments using the AdamW (Loshchilov & Hutter, 2019) optimizer, with a batch size of 1024. We linearly warmup the learning rate to $4e^{-3}$ in 20 epochs then lower it with a cosine scheulder to a final value of $1e^{-6}$. We use a layer-wise decay of 0.65, and a weight decay of 0.05. We do not use layer scale. We use 8 32GB V100 GPUs for our fine-tuning experiments.
@@
Linear-probing classification on ImageNet-1k. For our linear-probing experiments, we mostly follow DINO (Caron et al., 2021b) and DINOv2 (Oquab et al., 2024). We train for 100 epochs and a batch size of 8192 in all our experiments. We use the CLS token only from the last layer of the ViT encoder. Following Oquab et al. (2024), we instantiate multiple linear layers that all take the same encoder output, but are trained with different learning rates for efficiency. This way, the larger backbone inference is done once per iteration, while multiple inferences are done on the much lighter linear layers. We specifically span learning rate values of $\{1e^{-5},2e^{-5},5e^{-5},1e^{-4},2e^{-4},5e^{-4},1e^{-3},2e^{-3},5e^{-3},1e^{-2},2e^{-2},5e^{-2},1e^{-1}\},$ and we linearly scale them using the formula $lr\times batchsize/256$. We report the best performing value. We use a stochastic gradient descent optimizer with the momentum set to 0.9 and no weight decay. We use 8 32GB V100 GPUs for our linear-probing experiments.
@@
ADE20K semantic segmentation. In our semantic segmentation evaluations, we also use the iBOT (Zhou et al., 2022) pipeline, which follows BEiT (Bao et al., 2022). Specifically, we use UperNet (Xiao et al., 2018) from the implementation in mmsegmentaion (Contributors, 2020). We fine-tune for 160k iterations using a batch size of 16 and image size of 512 $\times$ 512. The AdamW optimizer is used with an initial learning rate of $8e^{-4}$ that is linearly warmed in the first 1500 epochs, then decays to 0 throughout training. We use a layer decay rate of 0.65 and weight decay of 0.05. We use Feature Pyramid Networks (FPNs) of four different scales to modify the feature map sizes generated by the ViT. Specifically, we upsample the output feature of the 4th block and 6th block by 4x and 2×, respectively, keep the output from the 8th block the same, and downsample the output feature of the 12th block by 2×. For data augmentation, we adopt the default settings in mmsegmentation, which includes random horizontal flipping, random re-scaling with a ratio range of $\left[0.5,2.0\right]$, and random photometric distortion. We use 4 32GB V100 GPUs for our semantic segmentation experiments.
@@
COCO object detection and instance segmentation. For object detection and instance segmentation, we use a Cascade Mask R-CNN (Cai & Vasconcelos, 2019; He et al., 2018) implementation with mmdetection (Chen et al., 2019), which produces both bounding boxes and instance masks. We follow the settings in iBOT (Zhou et al., 2022) in that we use multi-scale training with the shorter size between 480 and 800 while the longer size is not larger than 1333. We use a learning rate of $1e^{-4}$, a weight decay of 0.05, and fine-tune for 12 epochs with the learning rate decayed with a rate of 0.1 at epochs 9 and 11. We use a layer decay rate of 0.75 and a batch size of 16. We generate hierarchical feature maps by taking the outputs from layers 4, 6, 8, and 12, and passing them to two deconvolutions, one deconvolution, identity mapping, and max-pooling, respectively. We do not use multi-scale testing. We use 4 32GB V100 GPUs for our object detection experiments.
Zero-shot classification. For zero-shot classification, we follow the standard implementation (Radford et al., 2021; Mu et al., 2021) of encoding class labels in a description, such as "a photo of a {label}," and calualting the consine similarity between the text and image embeddings. We consider the highest generated similarity for each image-label pair to be the predicted class. For Harmony, we use the teacher text encoder.
Appendix B CC12M Zero-shot classification performance
We evaluate zero-shot classification performance of Harmony and CLIP, both trained on the CC12M dataset, across a diverse set of general vision benchmarks. As shown in Table 8, Harmony outperforms CLIP on average.
Table 8: Zero-shot evaluations on CC12M. We compare ViT-B Harmony against CLIP across a range of general and natural image benchmarks for zero-shot classification. Harmony continues to outperform CLIP on average.
  Average Cal101 STL-10 Kin700 MNIST CLEVR INet-A INet-O INet-R
CLIP 36.6 72.8 94.8 26.6 10.7 16.2 7.9 39.5 44.1
Harmony 39.6 76.0 96.7 30.5 9.6 12.7 12.2 36.0 54.2